D Question 1 1 pts You will save $1,000 per year You can earn 8% on your deposits. How much will you have 15 years from now At Round your answer to 2 decimals places, for example 100.21 Question 2 1 pts You will save some amount of money for 3 years. You can earn 2% on your savings. What is the FV factor for an annuity in this problem? Round your answer to 4 decimals, for example 10.1234. DI Question 3 1 pts You are planning to withdraw some amount of money every year for the next 23 years. You earn 6% on your deposits. What is the PV factor of an annuity for this problem? Round your answer to 4 decimals, for example 8.1234.
Answer #1

1. PMT = 1000, Rate = 8%, Time = 15 years
FV = 1000*((1+8%)15-1)/8% = 27152.11

2. FV factor = ((1+r)n-1)/r = ((1+2%)3-1)/2% = 3.06

3/ PV factor = (1-(1+r)-n)/r = (1-(1+6%)-23)/6% = 12.30

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well